OpenStudy (anonymous):

Solve the following quadratic equation using the quadratic formula and then choose the correct solution set. 6x^2 - 7x + 2 = 0 a) {-1/2,2/3} b) {1/2,2/3} c) {1/2,3/2}

OpenStudy (mertsj):

quadratic formula is: \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (mertsj):

In your problem, a = 6, b = -7, c = 2

hero (hero):

I don't know why they ask to solve by quad formula when it is already factorable. That sort of defeats the purpose.

OpenStudy (anonymous):

I keep getting weird answers

OpenStudy (anonymous):

is it b?

OpenStudy (mertsj):

Post your work and I will trouble shoot it.

OpenStudy (mertsj):

Yes. It is b

OpenStudy (mertsj):

Good job.

OpenStudy (anonymous):

aright thanks. sry I didn't know how to type my work out

OpenStudy (mertsj):

np

hero (hero):

You can also factor it 6x^2 - 7x + 2 = 0 6x^2 - 4x - 3x + 2 = 0 2x(3x - 2) - 1(3x - 2) = 0 (3x - 2)(2x - 1) = 0 3x - 2 = 0 2x - 1 = 0 x = 2/3 x = 1/2 For your purposes, however, quad formula might be easier for you.